OpenStudy (anonymous):

Solve the system algebraically. 4x - 5y = 3 x = 3 - y What is the value of x?

OpenStudy (anonymous):

Use substitution, put 3-y for x, 4(3-y) -5y = 3 to solve for y

OpenStudy (anonymous):

then when you have y put it in x=3-y to get x

OpenStudy (anonymous):

The problem is that I don't have a calculator. :/

OpenStudy (anonymous):

what do you need a calc for you can do 4(3-y) -5y =3 without a calc